OpenStudy (anonymous):

find the slope-intercept for the line satisfying the following conditions x intercept 2, y intercept 2/3.

OpenStudy (anonymous):

y = mx + b. This is slope-intercept form. As you know, to get the y intercept, we must set x = 0. So: y = m(0) + b y = b --> This is our y-intercept. Since we know that y-int is 2/3, we replace y with 2/3 and write it like this: 2/3 = b --> So we now know, that our b value in y = mx + b is 2/3. And as a fact, b in y = mx + b, is always the y-intercept. So now we re-write the equation with the info we have know and have verified: y = mx + 2/3 We also no that x-intercept is 2, and to get the x-intercept, we must set y as 0 since when a graph intersects the x-axis, the y-value there must be 0. After setting y as 0, we solve for x to get the x-int. So: 0 = mx + 2/3 -2/3 = mx (-2/3) / m = x --> We know that x-intercept is 2, so here we can plug in 2 for x. (-2/3) / m = 2 --> Now to get the equation for line, we only need the slope which is m, so now we solve for m. (-2/3) / 2 = m m = -1/3 --> Now we have m both the slope and y-intercept to put in y = mx + b. Thus the equation is: y = (-1/3)x + 2/3

OpenStudy (anonymous):

the slope intercept of the line would be?

OpenStudy (anonymous):

Are you asking for the slope and y-intercept or for the slope-intercept equation of the line?

OpenStudy (anonymous):

the slope intercept form of the line